OpenStudy (anonymous):

Arrange the steps to solve the equation x+3−−−−−√−2x−1−−−−−√=-2.

OpenStudy (anonymous):

\[\sqrt{x-3} -\sqrt{2x-1} =-2\]

OpenStudy (anonymous):

Arrange the steps to solve the equation. HELPP!!!

OpenStudy (anonymous):

square both sides first

OpenStudy (anonymous):

this will give you (x-3)-(2x-1)=4

OpenStudy (anonymous):

can you solve it from here?

OpenStudy (anonymous):

i need step by step

OpenStudy (anonymous):

ok, did you understand the first step though

OpenStudy (anonymous):